Triangles HIJ and MNO are similar. The length of the sides of HIJ are 111, 96, and 108 centimeters. The length of the smallest side of MNO is 288 centimeters, what is the length of the longest side on MNO?
manual responses are 315 cm , 324 cm, 333cm 945cm. I don't even know where to start with only one side of MNO given.
: Triangles HIJ and MNO are similar. The length of the sides of HIJ are 111, 96, and 108 centimeters. The length of the smallest side of MNO is 288 centimeters, what is the length of the longest side on MNO?
manual responses are 315 cm , 324 cm, 333cm 945cm. I don't even know where to start with only one side of MNO given.

Answer by scott8148(2719)

the triangles are similar, which means that the ratio between corresponding sides is constant

smallest is to smallest as longest is to longest __ let x="longest side"

288/96=x/111 __ 3=x/111 __ 333=x